OpenStudy (anonymous):

Find a, b, and h so that f(x) = asin(b(x-h). round to two decimals if needed, f(x)= 2sinpix + 3cospix

OpenStudy (tkhunny):

Start with this: \(\sqrt{2^{2} + 3^{2}}\)

OpenStudy (anonymous):

4+9? 13.

OpenStudy (tkhunny):

You lost the square root. \(a = \sqrt{13}\)

OpenStudy (anonymous):

oh! thank you!!

OpenStudy (tkhunny):

See where the 2 and 3 came from?

OpenStudy (anonymous):

yeah, the 2sin and the 3cos

OpenStudy (tkhunny):

Yup. Coefficients of the two individual functions. This works ONLY if they have the same period and phase shift. Speaking ofr same "Period", the resulting function must have the same period, so \(b = \pi\).

OpenStudy (tkhunny):

Well, the lasst piece you'll need is atan(3/2). Good luck!